OpenStudy (anonymous):

_____ √x+10=x-2 x+10 is all together under √ symbol

OpenStudy (anonymous):

\[\sqrt{x+10}=x-2\] (Take note: For \(\sqrt{x+10}\) to be defined, you must have \(x\ge-10\).) Start off by squaring both sides: \[\left(\sqrt{x+10}\right)^2=(x-2)^2\\ x+10=(x-2)^2\\ x+10=x^2-4x+4\\ x^2-5x-6=0\]

OpenStudy (anonymous):

is the next line 0=-6x+1x

OpenStudy (anonymous):

then zero principle property?

OpenStudy (anonymous):

You're right about the zero product property, but not with what you wrote. I think you meant to factor: \[(x-6)(x+1)=0\]

OpenStudy (anonymous):

hey will you give me a problem to see if i get it on each step....only if im wrong you correct me

OpenStudy (anonymous):

but on that one the ultimate answer is x=1 and x=6?

OpenStudy (anonymous):

Close: x=-1 and x=6

OpenStudy (anonymous):

yes i need to pay closer attention to pos and neg

OpenStudy (anonymous):

!hey will you give me a problem to see if i get it on each step....only if im wrong you correct me

OpenStudy (anonymous):

Try this one \[\sqrt{x-1}=-(x+1)\]

OpenStudy (anonymous):

x-1=(x+1)^2

OpenStudy (anonymous):

Yes, so far so good

OpenStudy (anonymous):

x-1=x^2+2x-1

OpenStudy (anonymous):

+1 at the end of the right side

OpenStudy (anonymous):

x+1=0 and x+1=0?

OpenStudy (anonymous):

\[x-1=x^2+2x+1\\ x^2+x+2=0\] Actually, it turns out you can't factor this easily. Have you learned the quadratic equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hold one moment....let me try work this one out on paper

OpenStudy (anonymous):

dude you have a negative one on the outside of the right side of radical eq?

OpenStudy (anonymous):

im lost on this one

OpenStudy (anonymous):

can we try a diffrent one

OpenStudy (anonymous):

Ah, sorry I got ahead of myself. There is no (real) solution to the equation I gave you. Try this one instead: \[x-3=\sqrt{5x+11}\]

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

0=x^2-11x-20

OpenStudy (anonymous):

Wait, something's not right... I must have made a mistake while coming up with the equation...

OpenStudy (anonymous):

but my steps were right though werent they

OpenStudy (anonymous):

Yes they were. I'll figure out what went wrong on my end.

OpenStudy (anonymous):

no need

OpenStudy (anonymous):

instead come up with new one

OpenStudy (anonymous):

maybe sq rt4x+1is what u meant

OpenStudy (anonymous):

No, I had something else in mind. It was supposed to be \(x+3\) on the left side.

OpenStudy (anonymous):

Anyway, here's another: \[\sqrt{x+4}=x+2\]

OpenStudy (anonymous):

x+3= square rt of 4x+11

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x=0 and x=-3?

OpenStudy (anonymous):

Right, but check those solutions.

OpenStudy (anonymous):

Only one of them works.

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

\(x=-3\) isn't a real solution. Can you see why? \[\sqrt{-3+4}=-3+2~?\]

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

yes i see why

OpenStudy (anonymous):

lets try another

OpenStudy (anonymous):

\[x+6=\sqrt{-x}\] this one also involves a solution that doesn't quite work.

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

x^2+13x+36=0

OpenStudy (anonymous):

You can factor that

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

x=-9 only checks out to be true the other was -4 but is false

OpenStudy (anonymous):

Are you sure about that? \[-9+6=\sqrt{-(-9)}\\ -3=\sqrt9\\ -3\not=3\] \(x=-9\) is not a real solution. \[-4+6=\sqrt{-(-4)}\\ 2=\sqrt4\\ 2=2\] \(x=-4\) is a real solution.

OpenStudy (anonymous):

what about this problem